OpenStudy (anonymous):

Determine the value of r so that the line through (6, -2) and (4, r ) has a slope of -3

OpenStudy (anonymous):

\[(-2-r)\div(6-4) = -3\] solve for r

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

i think its r=4

OpenStudy (anonymous):

how did you get 4?

OpenStudy (anonymous):

(-2-r)=-3(6-4)

OpenStudy (anonymous):

so r=4?

OpenStudy (anonymous):

yea try to plot it